OpenStudy (lifeisadangerousgame):

I have to evaluate cos (11pi/4) and sin (11pi/4)

OpenStudy (lifeisadangerousgame):

I did it the way I thought was right and I ended up with (sqrt2/2, sqrt2,2) but it's cos needed to be negative. They said the reason was that 11pi/2 is 2pi + pi/2 + pi/4 but I don't really understand that.

OpenStudy (er.mohd.amir):

(1/2)*2*sin (11pi/4) cos (11*pi/4) =(1/2) sin 22pi/4

OpenStudy (lifeisadangerousgame):

I still don't understand

OpenStudy (er.mohd.amir):

2 sin x cos x=sin 2x use this first

OpenStudy (er.mohd.amir):

now sin 11pi/2=sin (5pi+pi/2)

OpenStudy (lifeisadangerousgame):

It still doesn't make sense to me, do I need to use a calculator to do this method? Because I can't on the exam

OpenStudy (er.mohd.amir):

sin (4pi+pi+pi/2)=sin(pi+pi/2)

OpenStudy (lifeisadangerousgame):

I'm honestly very lost

OpenStudy (er.mohd.amir):

2pi is period of sin x that is sin(2pi+x)=sin x

OpenStudy (lifeisadangerousgame):

Okay